OpenStudy (anonymous):

2 over x+1 = 5 over x+3 how do i solve this?

OpenStudy (anonymous):

\[\frac{2}{x+1}=\frac{5}{x+3}\iff 2(x+3)=5(x+1)\] you good from there?

OpenStudy (anonymous):

how did you make the nice fractions

OpenStudy (anonymous):

latex \frac{a}{b}

OpenStudy (anonymous):

\[2x+6=5x+5\] \[6=3x+5\] \[1=3x\] \[x=\frac{1}{3}\] if my algebra is correct

OpenStudy (anonymous):

so ho are you making those fractions...when i post a questions i couldn't find anything to do that, i just had to type it...is there special characters?

OpenStudy (anonymous):

but i see how u did it i get it

OpenStudy (anonymous):

i am using latex. there is a latex study group if you want to look at it

OpenStudy (anonymous):

where do i go for that

OpenStudy (anonymous):

it is here somewhere you are in the math study group. there are lots of others

OpenStudy (anonymous):

i see..

OpenStudy (anonymous):

thx for help.

OpenStudy (anonymous):

yw